Using method of integration find the area of the triangle region  formed  whose  vertices are A  (2,-2),B (4,3) andC (1,2)

Grade:12

1 Answers

Ramit
5 years ago
A(2,-2) B(4,3) C(1,2).    X1=2   X2=4  X3=1
Y1=-2  Y2=3  Y3=2
Area of triangle ABC=1/2[X1(Y2-Y1)+X2(Y3-Y1)X3(Y1-Y2)]
=1/2[2(3-2)+4(2-(-2))+1(-2-3)]
=1/2[(1)+4(4)+(-5)]
1/2[2+16-5]
=1/2[2+11]=13/2  and = 13/2